OpenStudy (blackops2luvr):

Identify sqrt(2) as either rational or irrational, and approximate to the tenths place.

OpenStudy (blackops2luvr):

@Jhannybean @Nnesha

OpenStudy (blackops2luvr):

Before you answer this question, am i correct for this one?: Solve x^2 = 121. =, ±11

OpenStudy (blackops2luvr):

Because 11^2 can be -11, 11 so its both ways.

OpenStudy (blackops2luvr):

Am i wrong?

OpenStudy (solomonzelman):

you are correct about \(\large\color{black}{ x^2=121 }\) that. \(\large\color{black}{ x=\pm11 }\)

OpenStudy (blackops2luvr):

OMG!!!!!! you have been gone for \(\small\sf\color{red}{SOO}\) long!

OpenStudy (blackops2luvr):

k

OpenStudy (solomonzelman):

the picture I posted about real numbers numbers should help with the question of the post;)

OpenStudy (solomonzelman):

I wrote "numbers" twice;) lol

OpenStudy (blackops2luvr):

XD okay ty.

OpenStudy (solomonzelman):

yes, so is \(\large\color{black}{ \sqrt{2} }\) a rational, or an irrational number?

OpenStudy (blackops2luvr):

Irrational

OpenStudy (solomonzelman):

yes:)

OpenStudy (solomonzelman):

and the approximation of \(\large\color{black}{\sqrt{2} }\) to the nearest 10th?

OpenStudy (blackops2luvr):

1.5?

OpenStudy (solomonzelman):

we know that: \(\large\color{black}{\sqrt{2} \approx 1.4142... }\) (symbol \(\large\color{black}{\approx }\) means "approximately equal)

OpenStudy (solomonzelman):

when you round a \(\large\color{black}{1 }\) does it become a \(\large\color{black}{10 }\) or a \(\large\color{black}{0}\) ?

OpenStudy (blackops2luvr):

0

OpenStudy (blackops2luvr):

Oh wait

OpenStudy (solomonzelman):

yes, so when you say \(\large\color{black}{\sqrt{2} \approx 1.41 }\) then is it \(\large\color{green}{\sqrt{2} \approx 1.5 }\) or \(\large\color{blue}{\sqrt{2} \approx 1.4 }\) ?

OpenStudy (blackops2luvr):

1.4

OpenStudy (solomonzelman):

yup, there you go!

OpenStudy (blackops2luvr):

ty

OpenStudy (solomonzelman):

np

OpenStudy (blackops2luvr):

i got one more

OpenStudy (solomonzelman):

sure

OpenStudy (blackops2luvr):

Ill tag you.

OpenStudy (solomonzelman):

okay;)